OpenStudy (anonymous):

a bucket is 15kg,when it is filled with waternupto3/5 of its total capacity and its weight is 19kg when it is filled upto 4/5 of its total capacity.find the weight of the bucket if it is filled completely with water.

OpenStudy (anonymous):

Solve the following for w and b where w = weight of the water and b is the weight of the bucket: {2/5 w + b = 15, 4/5 w + b = 19} w=10 and b=11 w+b=21kg